determine the length of the unknown side (c) of a triangle 10cm and 24cm and determine the perimeter of the triangle and determine the area of the triangle.
Answer by TheInstructor(29)

a^2+b^2 = c^2
10^2 + 24^2 = 676
squareroot of 676 = 26
Therefore c = 26cm

Perimeter: 10 + 24 + 26 = 60 cm
Area = 1/2 x 10 x 24 = 120cm^2

:)